# DOSA-MO: Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems improves biomarker discovery

### 27 Dec 2023 v1

@@ 2.2.1 Pseudocode formulation of the DOSA-MO algorithm
In order to wrap any MO optimizer, the DOSA-MO must be polymorphic, so we define an abstract class MultiObjectiveOptimizer representing a generic MO optimizer (Pseudocode 1). It has a single method optimize that works with provided objectives and training data. The DOSA-MO is a MultiObjectiveOptimizer itself (Pseudocode 2).
@@
The method new is just a simple constructor that saves the polymorphic parts of the algorithm: the tuningOptimizer, a MO optimizer used to create the samples for the adjusting regression, the actual regression model (adjusterLearner), and the mainOptimizer that is the MO optimizer that uses the objectives adjusted by the regressions to produce the results for the user.
The DosaMO implementation of optimize first organizes the data into folds. The resulting object, foldsData, contains the data itself and also the description of how it is partitioned into folds. DosaMO then executes the tuningOptimizer on each fold and collects the results, that are a set of solutions for each fold. The set of solutions returned by each optimizer is optimizer-dependent in general, but in our experiments we used the Pareto front of all the solutions that were explored. For each objective obj the algorithm assigns weights to the solutions: for each tuningOptimizer result set h, the solutions receive a weight that is proportional to the partial derivative of the HV of the belonging result set with respect to the solution and the current objective obj. An adjuster regression model is trained for the current objective obj using the function trainAdjuster that receives in input a regression model adjusterLearner, the current objective obj, the tuningOptimizer solution set for each fold, the data including folds information (foldsData), and the weigths of the samples (weights). The returned adjuster predicts how much the fitness of a solution changes between the training sample set and an unseen testing sample set. The function trainAdjuster has its own detailed description below. Using the previous objective obj and the adjuster, a new adjusted objective is created that when evaluating a solution first uses the old fitness function to compute a temporary fitness, then adjustes this fitness by subtracting the prediction of the adjuster. Finally, DosaMO runs the mainOptimizer on the whole trainingData, using the adjusted objectives instead of the original objectives.
The trainAdjuster function trains a fitness adjuster regression model for one of the objectives using as samples the solutions resulting from running the tuningOptimizer on all the folds defined by foldsData. The solutions are weighted: each solution has a weight proportional to the HV partial derivative with respect to the considered objective, with the weights for each fold that sum to 1.
@@
For each fold i as defined by foldsData, trainAdjuster prepares the samples for training the adjuster regressor (Pseudocode 3). The samples are prepared separately for each fold, then used together in training. The function evaluateWithCV assigns two fitnesses to each solution: one computed on the train data of the current fold i, and another computed on the test data. It also computes the standard deviation of the train fitness computed with the boostrap method. The differences between train and test performance are the values that the regression will learn to predict. In order to do that the regression has 3 input meta-features: the original fitness, i.e. the performance on the train data, the standard deviation of the original fitness, and the number of the features that are included in the solution (the number of genes that are included in the biomarker in our case study).
The DOSA-MO is a wrapper algorithm that uses an MO optimizer to produce the training data for the regression models and a (possibly different) MO optimizer that will run with the adjusted objectives to produce the final results for the user. In our case study we used NSGA3-CHS algorithm as both tuning and main optimizer. The parameters of population size and number of generations were lower for the tuning optimizer in order to reduce the computational load. Modified Non-Dominated Sorting Genetic Algorithm II (NSGA2) algorithms NSGA2-CH and NSGA2-CHS were first introduced in [5] and further validated in [6]. In this work we analogously modified NSGA3 with the same capabilities of NSGA2-CHS obtaining NSGA3-CHS. For this purpose we designed a more general algorithm: NSGA*.
class Nsga* inherits MultiObjectiveOptimizer
Ψ
   method new(
         nGenerations, popsize, featureImportance,
         sort, tournament, mutation,
         cloneRepurposing):
      self.nGenerations = nGenerations
      self.popsize = popsize
      self.featureImportance = featureImportance
      self.sort = sort
      self.tournament = tournament
      self.mutation = mutation
      self.cloneRepurposing = cloneRepurposing
   method optimize(objectives, trainingData):
      hof = createNewHallOfFame()
      featImp = self.featureImportance(trainingData)
      pop = createNewRandomIndividuals(
            featImp, self.popsize)
      for 1::self.nGenerations:
         pop = evaluate(pop, objectives)
         pop = self.sort(pop)
         offspring = self.tournament(pop)
         offspring = crossover(offspring)
         offspring = self.mutation(offspring)
         pop = concat(pop, offspring)
         if self.cloneRepurposing:
            pop = ReplaceClonesWithNewIndividuals(
                  pop, featImp)
         pop = evaluate(pop, objectives)
         hof = updateHallOfFame(hof, pop)
         pop = self.sort(pop)
         pop = pop[1..self.popsize]
      return hof
Pseudocode 4 NSGA* class definition.
Pseudocode 4 describes the algorithm NSGA*, a further generalization of the generalized Nsga2 algorithm presented in [5]. NSGA* generalizes NSGA2 and NSGA3, together with their modifications NSGA2-CH, NSGA2-CHS, NSGA3-CH, and NSGA3-CHS. The differences between the algorithms derived from NSGA2 and NSGA3 reside in the tournament and in the sort routines. NSGA2 uses a tournament by position: the individuals are paired if their position divided by 2 is the same [16]. NSGA3 uses a random tournament where pairs are selected randomly with replacement [15]. Both NSGA2 and NSGA3 use a hyerarchical sort where individuals are first sorted by their domination front. The secondary sorting is different: it is based on the crowding distance in NSGA2 [16] and on reference point niching in NSGA3 [15]. The modifications CH and CHS use a primary sorting according to the clone index ([5]), and the original sorting of NSGA2 or NSGA3 as secondary sorting.
The class Nsga* has a constructor that memorizes the algorithm-specific parameters and strategies: the number of generations, population size, feature importance function, sorting algorithm, tournament strategy, mutation operator, and flag for the use of clone repurposing. Since the DOSA-MO uses adjusted objective functions, NSGA* takes an objectives object in input, it is used by the evaluate function and represents the NSGA* ability to run with different objective functions. The algorithm is similar to generalized NSGA2 [5], but with the added possibility of personalizing the tournament strategy, sorting algorithm and objectives. It can be used as inner MO optimizer by the DOSA-MO and can be specialized also as NSGA3, NSGA3-CH, and NSGA3-CHS.
2.2.2 Considered adjusting regression models
Five regression models have been considered in the experimental validation for adjusting the fitness functions. All of them use the sample weights, computed by the DOSA-MO as the partial derivatives of the HV as explained above.
Dummy.
The simplest regression model learns the weighted median.
ptree.
The pruned decision tree regression model, minimizing the weighted absolute error. The tree is pruned with the Minimal Cost-Complexity Pruning algorithm [17]. The complexity parameter for the pruning is optimized by running a 5-fold CV on its training data.
RFReg.
The random forest regression, minimizing the weighted absolute error.
SVR.
The epsilon-support vector regression with Gaussian kernel type [18]. It uses an l2 regularization penalty. We use the default parameters of $C=1$ and $\epsilon{}=0.1$.
rSVR.
The randomized SVR uses random search with 5-fold CV on its training data to optimize the parameters $C$ and $\epsilon{}$.
We include in the tests the unadjusted NSGA3-CHS, that is equivalent to use regression models that predict always 0 (shortened as “zero” in the following).
2.2.3 Limit the computational overhead for adjustments
Running the tuning optimizer in a nested k-fold CV in order to generate the samples used for training the regression models imposes a computational overhead with respect to the cost of running the main optimizer without any adjustment to the fitness functions. An high overhead could make the DOSA-MO impractical in real cases. In our experimental validation, that uses GAs as inner MO optimizers, we limited the computational overhead by using in the tuning optimizer a smaller population and less generations than in the main optimizer.
The simplifying assumption is made for the computational cost of the GAs to be proportional to the number of individual evaluations multiplied by the number of samples used for evaluating the individuals (in training algorithms the cost is typically at least linear in the number of samples, since the algorithm has to at least iterate through them), and for the number of evaluations to be in turn proportional to the population size multiplied by the number of generations (this is in fact an upper bound considering that individuals equal to previously evaluated ones might not need to be evaluated again). We define a parameter $\mu{}$ as the ratio between the number of individuals evaluated during the tuning phase with respect to the main optimization phase.
In order to have a computation time of the DOSA-MO approximately double than the time required by the unadjusted optimizer, $\mu{}$ is set to 1 in our experiments.
In the case of internal-external CV, let $\eta{}$ be the number of external folds. We use the same number of folds also for the generation of the training samples for the regression models, so the tuning optimizer is executed that number of times for each execution of the DOSA-MO. Let $\rho{}$ be the population size for the main optimizer, and $\rho{}^{\prime}$ the population size for the tuning optimizer. We compute $\rho{}^{\prime}$ with the following equation.
$\rho{}^{\prime}=\left\lfloor\rho{}\sqrt{\frac{\mu{}}{\eta{}-1}}\right\rceil$ (3)
[…]
Let $\gamma{}$ be the number of generations for the main optimizer, and $\gamma{}^{\prime}$ the number of generations for the tuning optimizer. Similarly, we compute $\gamma{}^{\prime}$ with the following equation.
$\gamma{}^{\prime}=\left\lfloor\gamma{}\sqrt{\frac{\mu{}}{\eta{}-1}}\right\rceil$ (4)
According to the previous assumptions, the cost of running the main optimizer $c$ is $c=\rho{}\gamma{}m$, with $m$ being the number of samples. The cost of running all the iterations of the tuning optimizer $c^{\prime}$ is the following.
$c^{\prime}=\eta{}\rho{}^{\prime}\gamma{}^{\prime}\frac{m(\eta{}-1)}{\eta{}}=\left\lfloor\rho{}\sqrt{\frac{\mu{}}{\eta{}-1}}\right\rceil\left\lfloor\gamma{}\sqrt{\frac{\mu{}}{\eta{}-1}}\right\rceil m(\eta{}-1)$ (5)
Ignoring the round operations that have just a small contribution it is easy to verify the desired ratio.
$c^{\prime}\approx\rho{}\gamma{}\mu{}m=mc$ (6)
So the computational cost of using the adjusted optimizer wrapper, including the main optimizer, is approximately equal to $(1+m)c$, or even lower if the cost of evaluating individuals grows more than linearly in the number of samples.
The external validation is faster since it does not have an outer k-fold CV. For the external validations, we have arbitrarily set $\eta{}=5$, resulting in samples for regression training to be acquired from 5 folds.
2.2.4 Computing the fitness functions variance
One of the high-level features of the solutions, used for prediction by the adjustment regression models, is the standard deviation of the original fitness.
Collecting the fitnesses of an individual (feature set in our case study) on the different folds and computing the standard deviation of a performance metric (for example, the balanced accuracy or the concordance index), would have a limited precision because of the number of the folds, that is necessarily contained to have an acceptable computational time. On the other hand, increasing the fold count would heighten computational demands, as the inner models need to be trained for each fold. Additionally, each fold fitnesses would be estimated on a smaller left-out sample set. Moreover, performing repeated CV would increase the number of evaluations but the same test samples would be reused.
Our method to calculate each solution’s fitness variances addresses these issues by performing bootstrap analysis within each fold of the k-fold CV, followed by aggregation of the results across all folds. The following steps describe how one of the fitnesses for one of the individuals is computed.
  1. 1.
    For each of the folds there are training and testing samples. The inner model is optimized on the training data. The fitness is computed on the whole testing data, then by using bootstrap on the testing data, the variance for the fitness is computed.
  2. 2.
    The results from each of the folds are aggregated considering the folds as strata in a stratified bootstrap. The variances in the different folds are assumed uncorrelated and combined with the equation for the variance of the mean of uncorrelated random variables: the variance of the mean is the sum of the variances divided by the square of the number of folds.
When cross-validating, there are two sources of performance variance: the composition of the training set affects the training process thus can lead to different predictors, and indirectly to different expected performance, while the composition of the test set directly impacts the expected performance [19]. It is known that there is no unbiased estimator of the variance of k-fold CV [20]. The described procedure accounts for the variance explained by the limited number of test samples, but only partly for the variance explained by the limited number of training samples, since the training samples are reused in different folds (this is a well known unavoidable limitation in k-fold CV [21]) and the bootstrap is applied on the test sets but not on the training sets, to avoid incurring in unfeasible computational costs. Despite the limitations, this estimation of the fitness variance appeared predictive of the overestimation in our experiments, thus justifying its inclusion in the high level features for the overestimation prediction.
2.3 Benchmarking datasets: description and pre-processing
We conducted benchmark studies with two primary goals: firstly, to identify gene expression-based biomarker sets for classifying cancer subtypes, and secondly, to determine similar biomarker sets for survival prediction in kidney and breast cancer patients.
Both case studies were addressed by using internal-external CV on TCGA [13] data. Additionally, the breast cancer case study includes an external validation with training on TCGA data and testing on SCAN-B data [14]. The latter test is particularly important as it demonstrates that the proposed algorithm can also be utilized by training predictive models for both classification and overestimation within a specific cohort (e.g. TCGA) and applying them in external cohorts (e.g. SCAN-B).
TCGA transcriptomic datasets were downloaded with the curatedTCGAData R-package version 2.0.1 from assays of type RNASeq2GeneNorm [22]. The retrieved data consists of upper-quartile-normalized TPM values. They were log transformed by applying $log_{2}(x+1)$.
The external gene expression-based transcriptomic dataset for breast cancer was obtained from the Gene Expression Omnibus (GEO) database (GSE96058) collected from the SCAN-B consortium. It includes FPKM log-transformed gene expression profiles. This data was already log-transformed, and we did not apply our own log-transformation to it. We applied for each value $x$ the function $max(x,0)$, then we excluded the genes with less than 30% non-zero values.
For our study, we utilized TCGA datasets specific to cancer types. For the breast cancer case study, we used the TCGA-BRCA dataset. Additionally, for kidney cancer, we considered a compendium of TCGA datasets, which includes: KICH (Kidney Chromophobe), KIRC (Kidney Renal Clear Cell Carcinoma), and KIRP (Kidney Renal Papillary Cell Carcinoma).
Each transcriptomic dataset was filtered by removing genes with zero variance and gene expression values were standardized before use. For both TCGA and SCAN-B, cancer patient samples were categorized based on the PAM50 cancer subtype signature, which is used to determine specific molecular subtypes of breast cancer. The subtypes identified by PAM50 that were used for our experiments are: Luminal A (LumA), Luminal B (LumB), HER2-enriched (Her2), Basal-like (Basal, which is often referred to as triple-negative) and normal-like (Normal) cancer. For the task of classifying cancer subtypes in TCGA kidney cancer patients, we focused on distinguishing clear-cell renal cell carcinoma (ccRCC), chromophobe renal cell carcinoma (ChRCC), and papillary renal cell carcinoma, which was further divided into two subtypes based on recent studies identifying distinct clinical categories. These are referred to as PRCC T1 and PRCC T2. Additionally, we included samples of normal non-cancerous tissues. This classification system is based on a study published by Ricketts et al. [23]. Overall survival data for TCGA kidney cancer is from Liu et al. [24].
Supplementary Fig. S1 shows the first two principal components (after standardization) and the number of mRNA-based profiles for each kidney cancer subtype. Supplementary Fig. S2 shows the same information for the samples that are used when the prediction of the overall survival is an objective. In this cases the normal tissues are not used because in TCGA they are collected from healty areas of cancer patients. Additionally, two samples are excluded because they are missing the survival information. Supplementary Fig. S3 shows the distribution of the survival outcomes in time.
Supplementary Fig. S4 shows the first two principal components (after standardization) and the number of mRNA-based profiles for each breast cancer subtype in TCGA data. The same information for the SCAN-B dataset is shown in Supplementary Fig. S5.
2.4 Experimental setup
Our tests include three datasets: TCGA kidney used for internal-external CV, TCGA breast used for internal-external CV and for the training part of external validation, and SCAN-B used for the testing part of external validation. The datasets and their preprocessing are described in Section 2.3.
Our MO problems include three objectives that measure the ability to classify the cancer subtypes, the ability to predict the overall survival, and the feature set size. These objectives are measured in the $[0,1]$ interval as per assumption of the algorithms. Subtype classification is measured with the balanced accuracy, survival prediction with the c-index, while the feature set size with the novel metric root-leanness. It was measured with leanness in a previous study [5]. The leanness is a value between 0 and 1 with higher values for solutions with less features. It is defined as $1/(n+1)$, with $n$ being the number of features. The leanness decreases sharply as the feature set size increases; consequently, the CHV is strongly impacted by the accuracy of the smallest biomarkers that use 1 or 2 features. Taking this into account, Cattelani et al. [6] provided an additional evaluation of the same solution sets by using a different measure for the impact of the number of features: $1/(\sqrt{n}+1$). This soft-leanness decreases less sharply as the set size increases. the root-leanness is even smoother that the soft-leanness and is defined as the root of the leanness: $\sqrt{1/(n+1)}$.
The program performs a 3-fold CV repeated 3 times for internal-external CV on the kidney dataset, while, for uniformity with previous works [5, 6], a 5-fold CV on the breast dataset. The MO optimizer used is the DOSA-MO, wrapping NSGA3-CHS as both tuning optimizer and main optimizer (Section 2.2.1). The hyperparameters of NSGA3-CHS used as main algorithm are population 500, generations 500, and 3 folds used for the evaluation of the individuals. The hyperparameters related to the tuning phase of the DOSA-MO are set as explained in Section 2.2.3. Each test is repeated 6 times, one for each of the 5 regressors described in Section 2.2.2, and one for the zero regressor (equivalent to the unadjusted NSGA3-CHS).
In this study, for the task of cancer subtype classification, NB and SVM were selected as the inner models. Meanwhile, Cox was employed as the inner model for the task of cancer survival prediction. The SVM uses the Radial basis function kernel, balanced class weight, and l2 regularization parameter $C=1$.
@@ 2.4 Experimental setup
The considered experimental setups are listed in Table 1.
In each k-fold CV execution, samples are stratified based on the included objectives. For balanced accuracy, stratification is by cancer subtype, and for overall survival, by survival status and time (binned into high or low for evenly distributed death events). If both these objectives are present, stratification combines both criteria.
During each GA call, for classification objectives, features not passing an ANOVA test (p-value 0.05) are removed from the current training samples. For survival objectives, features that fail a Wald test (p-value 0.05) in a Cox regression are discarded. If both objectives apply, features failing both tests are dropped.
For each combination of validation type, datasets, objectives, adjusting regression model, and classification inner model, we calculated the MO performance error and Pareto delta to assess the difference between expected and actual test sample performance. The CHV was also computed as an overall performance indicator for the solution sets. To the best of our knowledge it is the only proposed generalization of both HV to CV scenarios and of single-objective CV to MO [6]. CHV takes into account the differences between the performance expected by the optimization algorithm and measured on the test samples and preserves the HV appreciated properties, in particoular the strict monotonicity [25]: if a set of solutions is strictly better, the CHV is guaranteed to be higher.

### 29 Feb 2024 v3

@@ 2.2 Pseudocode formulation of the DOSA-MO algorithm
In order to wrap any MO optimizers, the DOSA-MO must be polymorphic, so we define an abstract class MultiObjectiveOptimizer representing a generic MO optimizer (Pseudocode 1). It has a single method optimize that works with provided objectives and training data. The DOSA-MO is a MultiObjectiveOptimizer itself (Pseudocode 2).
@@
The method new is just a simple constructor that saves the polymorphic parts of the algorithm: the tuningOptimizer, a MO optimizer used to create the samples for the adjusting regression, the actual regression model (adjusterLearner), and the mainOptimizer that is the MO optimizer that uses the adjusted objectives to produce the results for the user.
The DosaMO implementation of optimize first organizes the data into folds. The resulting object, foldsData, contains the data itself and also the description of how it is partitioned into folds. DosaMO then executes the tuningOptimizer on each fold and collects the results: a set of solutions for each fold. The set of solutions returned is optimizer-dependent in general, but in our experiments we used the non-dominated front of all the solutions that were explored. For each objective obj the algorithm assigns weights to the solutions: for each tuningOptimizer result set h, the solutions receive a weight that is proportional to the partial derivative of the HV of the belonging result set with respect to the solution and the current objective obj. An adjuster regression model is trained for the current objective obj using the function trainAdjuster that receives in input a regression model adjusterLearner, the current objective obj, the tuningOptimizer solution set for each fold, the data including folds information (foldsData), and the weigths of the samples (weights). The returned adjuster predicts how much the fitness of a solution changes between the training sample set and an unseen testing sample set. The function trainAdjuster has its own description below. Using the previous objective obj and the adjuster, a new adjusted objective is created that when evaluating a solution first uses the old fitness function to compute a temporary fitness and its SD, then adjustes this fitness by subtracting the prediction of the adjuster. Finally, DosaMO runs the mainOptimizer on the whole trainingData, using the adjusted objectives instead of the original objectives.
The trainAdjuster function trains a fitness adjuster regression model for one of the objectives using as samples the solutions resulting from running the tuningOptimizer on all the folds defined by foldsData. Each solution is assigned a weight proportional to the HV partial derivative with respect to the considered objective, with the weights for each fold that sum to 1.
@@
For each fold i as defined by foldsData, trainAdjuster prepares the samples for training the adjuster regressor (Pseudocode 3). The samples are prepared separately for each fold, then used together in training. The function evaluateWithCV assigns two fitnesses to each solution: one previously computed on the train data of the current fold i (using a nested k-fold CV in our case study), and another computed on the test data. It also computes the SD of the train fitness with the bootstrap method (Supplementary Section 1.4). The differences between train and test performance are the values that the regression will learn to predict. The regression has 3 input meta-features: the original fitness, i.e. the performance on the train data, the SD of the original fitness, and the number of the features that are included in the solution (number of genes in our case study).
2.3 Case study overview
We have benchmarked DOSA-MO in the context of MOFS for cancer biomarker discovery. Our goal was to identify biomarkers for classifying cancer subtypes and survival prediction in kidney and breast cancer patients. This was done using gene-based transcriptomic datasets from The Cancer Genome Atlas (TCGA) project [18]. For breast cancer, an additional cohort from The Sweden Cancerome Analysis Network - Breast (SCAN-B) [19] serves as an external validation set. A description of the datasets and their preprocessing is in Supplementary Section 1.5. As wrapped MO optimizer, we used a novel modification of NSGA3 [15]: NSGA3-CHS, with NB or Support Vector Machine (SVM) as inner classifiers, and Cox as inner survival model. We used both internal k-fold CV and external validation to compare the unadjusted optimizer (abbreviated as “zero”) with adjustments by 5 different regression models: weighted median (dummy), pruned decision tree (ptree), random forest (RFReg), support vector regression (SVR), and SVR with optimized regularization parameters (rSVR). The regression models are described in detail in Supplementary Section 1.2.
@@ 2.3 Case study overview
NSGA3-CHS is an instantiation of the more general algorithm NSGA*. They are both defined in Supplementary Section 1.1. The considered experimental setups are listed in Table 1 and described in detail in Supplementary Section 1.6, where it is also defined the root-leanness, used as fitness function for the parsimony of the feature sets. Classification and survival prediction performance are measured respectively with balanced accuracy and concordance index (c-index).
2.4 Measuring overestimation in multi-objective problems
We propose two new methods to measure the error of the performance estimates for the solutions to MO problems: MOPE and $P_{\Delta}$. The data-driven approach complicates the measurements, as CV yields varied performance metrics. These include the performance anticipated by the optimizer, based on training data, and the performance measured on left-out/new samples post-optimization. To our knowledge, there is no established metric for the error in evaluating the predictive performance of solutions in MO problems within a CV framework. In SO scenarios, one might simply assess the absolute difference between the fitness expected from training data and the fitness observed on new data. However, in MO scenarios, it is crucial to consider each solution’s contribution to the Pareto front. We introduce two novel metrics to measure this estimation error in MO CV setups.
2.4.1 Multi-objective performance error (MOPE)
We define the MOPE $E_{\upsilon}$ starting from the HV [17] computed on the train performance ($H_{\iota}$) and the Cross HyperVolume (CHV) ($H_{\upsilon}$). $H_{\iota}$ and $H_{\upsilon}$ have been formally defined by Cattelani et al. [8]. Since $H_{\upsilon}$ is a family of functions, with the specific instantiation that depends on the user provided function $\upsilon$, $E_{\upsilon}$ is a family of functions too.
Where $X$ encodes the performance of the solutions evaluated on the train data, and $X^{\prime}$ the performance evaluated on the test data. In our case study we use the same instantiation of $\upsilon$ as in Cattelani et al.: the function $\lambda$ [8].
2.4.2 Pareto delta ($P_{\Delta}$)
The MOPE, derived directly from the training HV and CHV, does not specify the exact sources of discrepancies between these measures. Although CHV accounts for differences between training and testing performance, variations in solution performance of opposite sign can offset each other, resulting in a lower MOPE despite significant train-test discrepancies. To tackle this issue, we introduce a supplementary metric, the $P_{\Delta}$. It has the property of being equal to 00 only when there is no difference in performance between train and test data for all elements of the solution set. To calculate the $P_{\Delta}$, we sum up the absolute error in fitness estimation for each solution, multiplying it by the derivative of the HV concerning that specific objective and solution. Then, we compute the average across all objectives.
Let $n$ be the number of solutions and $m$ the number of objectives. Let $X$ be an $n\times m$ matrix where $x_{i,j}$ is the train performance for the $j^{th}$ objective of the $i^{th}$ solution, with $1\leq i\leq n$ and $1\leq j\leq m$. Similarly, let $X^{\prime}$ be an $n\times m$ matrix where $x^{\prime}_{i,j}$ is the test performance for the $j^{th}$ objective of the $i^{th}$ solution. Let $H_{\iota}$ be the HV computed on the train performance $X$. $\partial H_{\iota}/{\partial x_{i,j}}$ is the partial derivative of the HV with respect to $x_{i,j}$. We define the $P_{\Delta}$ function as